How would you solve for "x" if...

ln(x^2-16)= 2ln(3)

Answer by ewatrrr(24785)

 
Hi,
ln(x^2-16)= 2ln(3)
ln(x^2-16)-ln(3^2)= 0
ln(x^2-16)/9) = 0
(x^2-16)/9 = e^0 = 1
x^2-16 = 9
x^2 = 25
x = ± 5
both Solutions check